There is a bag filled with 3 blue,2 red and 4 green marbles.A marble is taken at random from the bag,the colour is noted and then replaced.Another marble is taken at random.What is the probability of getting 2 different colours?


1
Expert's answer
2022-02-04T08:19:36-0500

Let A - "two different"

"P(A)=1-P(\neg A)"

"\neg A" - "both same"

Let B - "both blue", C - "both red", D - "both green", then

"P(\neg A)=P(B)+P(C)+P(D)=({\frac 3 {3+2+4}})^2+({\frac 2 {3+2+4}})^2+({\frac 4 {3+2+4}})^2={\frac{29} {81}}"

So, "P(A)=1-{\frac {29} {81}}={\frac {52} {81}}"